EMS Council Meeting

The Council of EMS meets on 20–21 July, 1996, in Budapest, at the the Mathematical Research Institute of the Hungarian Academy of Sciences, district 5, Realtanoda utca 13-15, Lecture room, beginning 9.00 a.m. The Agenda of the meeting has been mailed to delegates in April and futher material will be sent in early June. Delegates who have not received the Agenda are kindly requested to contact the Secretariat (e-mail: makelainen@cc.helsinki.fi fax: +358-0-19123213).